OpenStudy (anonymous):

which is the equations for the line perpendicular to y=5/3x+11 1/3 and containing p(-2,3) y-2=-3/5(x-3) y=-5/3x+4 1/3 y=-3/5x+4 1/5 y=3/5x+4 1/5

OpenStudy (anonymous):

basic information perpendicular lines have slopes that are negative inverses of each other if one has the slope and a point of that slope:point-slop form can be used to find the line's equation

OpenStudy (anonymous):

point-slope form:y-y1=m(x-x1)

OpenStudy (anonymous):

so the last one?

OpenStudy (anonymous):

well the slope would be -3/5

OpenStudy (anonymous):

it is not B, or D

OpenStudy (anonymous):

but they already have a -5/3 in the first part of the question

OpenStudy (anonymous):

? y=5/3x+11 1/3 <---did you make a typo?

OpenStudy (anonymous):

oh yes my bad. its supposed to be y=-5/3x+11 1/3

OpenStudy (anonymous):

ok then the slope would be 3/5 so yes D

OpenStudy (anonymous):

yay thank you!